# Predicting the adhesion and delamination strength of carbon films on metals by high-throughput ab initio calculations

### v1

@@ III.5 Charge density accumulation at the interface
The structural modifications of the diamond reconstructions directly affect the electronic charge redistribution at the interface. Previous studies have shown that the redistribution of electronic charge upon contact, quantified by the descriptor (Eq.2), correlates well with adhesion across a wide variety of interfaces, from weakly interacting Van der Waals systems (e.g., graphene) to covalently bonded materials such as diamond [59], as well as heterogeneous metallic junctions [46].
In the present case, the semiconductor–metal interaction introduces additional complexity, as surface carbon chains can undergo rehybridization and other structural rearrangements, modifying the local electronic environment. Despite this, a strong correlation is still observed between adhesion energy and , with a value of 71 (Fig.9), confirming that remains a reliable predictor even for this more intricate class of interfaces. Deviations from the linear trend arise from the interplay of chemical specificity, structural rearrangements and local orbital interactions with the metal.
IV CONCLUSIONS
@@
In this work, a high-throughput ab initio investigation of diamond/metal interfaces has been carried out, systematically analyzing the adhesion of 60 combinations involving a broad range of metallic substrates and four low-index diamond terminations (C(111), C(111)Pd, C(110), C(100)), representative of the different types of bonds in DLC coatings. The study employs an automated workflow implemented within the TribChem software, which ensures a consistent and reproducible protocol for the construction and characterization of solid interfaces. In particular, the interface supercells were built using the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while minimizing both lattice mismatch and angular distortion. All interfaces were constructed with a metallic strain below 3, ensuring that the calculated adhesion energies are not affected by unphysical distortions.
@@
The results show that adhesion is primarily governed by the intrinsic reactivity of the metallic component and by the type of carbon hybridization present at the surface of the diamond. Carbide-forming metals such as Cr, Ti, W and V exhibit the strongest adhesion, in agreement with their experimentally recognized role as interlayer materials that enhance DLC adhesion to metallic substrates. In contrast, noble and weakly reactive metals (Ag, Au, Cu, Zn, Mg) display poor adhesion due to their filled valence shells and low surface energies. Among the diamond terminations, the C(111) surface with sp3 bonding shows the highest adhesion, whereas surface graphitization, as observed in C(111)Pd, considerably weakens interfacial bonding.
@@
By comparing the work of separation () of the diamond/metal interface with the cohesive energy of the corresponding bulk metal, it is possible to predict the likely fracture location under tensile load. Interfaces involving C(111)Pd consistently exhibit fracture at the interface, reflecting the weak bonding due to sp2 terminations. For metals with low cohesive energies, such as Mg, Zn and Al, fracture can occur within the metal bulk when in contact with highly reactive diamond surfaces, indicating that the interface is stronger than the metallic cohesion. Conversely, transition metals with partially filled d-shells (e.g., W, Cr, Mo, Fe) exhibit high bulk cohesion, preventing failure from taking place within the metal phase.
Analysis of the radial distribution function provides insights on the structural response of diamond surface reconstructions upon contact with metals. In C(111)Pd, the peak at approximately 1.44 Å, corresponding to the zig-zag surface chains, is largely preserved in the presence of weakly interacting metals such as Au and Ag, indicating minimal perturbation of the sp2 network. In contrast, Cr, Al, Cu and Pt induce significant dimerization and partial rehybridization of the surface sp2 bonds. In general, reactive metals like W, Cr, Mo, Fe, V and Ti are able to reduce the number of surface sp2 bonds by up to 50. Hence, it can be concluded that strong perturbation and reduction in the number of surface sp2 bonds in diamond coatings directly enhance adhesion, explaining why these metals are commonly used as interlayers to improve bonding between DLC films and metallic substrates.
The structural modifications are mirrored in an electronic rearrangement at the interface. Volumetric charge density differences show that metals inducing stronger structural perturbations promote more pronounced charge rearrangement, while weaker interacting metals cause only minor redistribution. Quantitatively, the charge accumulation descriptor, , remains strongly correlated with adhesion, confirming that electronic charge redistribution upon interface formation is a good predictor of adhesion strength, even in the presence of semiconductor-metal bonding.
@@ Acknowledgements.
These results are part of the ”Advancing Solid Interface and Lubricants by First Principles Material Design (SLIDE)” project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 865633). The authors acknowledge funding by the European Union - NextGenerationEU (National Sustainable Mobility Center CN00000023, Italian Ministry of University and Research Decree n. 1033 - 17/06/2022, Spoke 11 - Innovative Materials Lightweighting) and by the European Union - NextGenerationEU (Spoke 6 - Mulstiscale Modelling Engineering Applications). The opinions expressed are those of the authors only and should not be considered as representative of the European Union or the European Commission’s official position. Neither the European Union nor the European Commission can be held responsible for them. Computational resources were provided by CINECA under the ISCRA initiative.References
- Maguire et al. [2005] P. Maguire, J. McLaughlin, T. Okpalugo, P. Lemoine, P. Papakonstantinou, E. McAdams, M. Needham, A. Ogwu, M. Ball, and G. Abbas, Mechanical stability, corrosion performance and bioresponse of amorphous diamond-like carbon for medical stents and guidewires, Diamond and Related Materials 14, 1277 (2005).
- Dearnaley and Arps [2005] G. Dearnaley and J. H. Arps, Biomedical applications of diamond-like carbon (dlc) coatings: A review, Surface and Coatings Technology 200, 2518 (2005).
- Hauert [2003] R. Hauert, A review of modified dlc coatings for biological applications, Diamond and Related Materials 12, 583 (2003).
- Wang et al. [2020] T. Wang, L. Huang, Y. Liu, X. Li, C. Liu, S. Handschuh-Wang, Y. Xu, Y. Zhao, and Y. Tang, Robust biomimetic hierarchical diamond architecture with a self-cleaning, antibacterial, and antibiofouling surface, ACS Applied Materials & Interfaces 12, 24432 (2020).
- Robertson [2002] J. Robertson, Diamond-like amorphous carbon, Materials Science and Engineering: R: Reports 37, 129 (2002).
- Lepicka et al. [2017] M. Lepicka, M. Gradzka-Dahlke, D. Pieniak, K. Pasierbiewicz, and A. Niewczas, Effect of mechanical properties of substrate and coating on wear performance of tin- or dlc-coated 316lvm stainless steel, Wear 382-383, 62 (2017).
- Scharf and Prasad [2012] T. W. Scharf and S. V. Prasad, Solid lubricants: a review, Journal of Materials Science 48, 511 (2012).
@@
- Auciello et al. [2004] O. Auciello, J. Birrell, J. A. Carlisle, J. E. Gerbi, X. Xiao, B. Peng, and H. D. Espinosa, Materials science and fabrication processes for a new mems technology based on ultrananocrystalline diamond thin films, Journal of Physics: Condensed Matter 16, R539 (2004).
- Love et al. [2013] C. Love, R. Cook, T. Harvey, P. Dearnley, and R. Wood, Diamond like carbon coatings for potential application in biological implants—a review, Tribology International 63, 141 (2013).
- Azzi et al. [2009] M. Azzi, M. Paquette, J. Szpunar, J. Klemberg-Sapieha, and L. Martinu, Tribocorrosion behaviour of dlc-coated 316l stainless steel, Wear 267, 860 (2009).
- Wang et al. [2002] R. Wang, C. Mercer, A. Evans, C. Cooper, and H. Yoon, Delamination and spalling of diamond-like-carbon tribological surfaces, Diamond and Related Materials 11, 1797 (2002).
- Lawes et al. [2007] S. D. A. Lawes, M. E. Fitzpatrick, and S. V. Hainsworth, Evaluation of the tribological properties of dlc for engine applications, Journal of Physics D: Applied Physics 40, 5427 (2007).
- Owens et al. [2015] A. G. Owens, S. Brühl, S. Simison, C. Forsich, and D. Heim, Comparison of tribological properties of stainless steel with hard and soft dlc coatings, Procedia Materials Science 9, 246 (2015).
- Luo et al. [2007] J. K. Luo, Y. Q. Fu, H. R. Le, J. A. Williams, S. M. Spearing, and W. I. Milne, Diamond and diamond-like carbon mems, Journal of Micromechanics and Microengineering 17, S147 (2007).
- Kasiorowski et al. [2020] T. Kasiorowski, J. Lin, P. Soares, C. Lepienski, C. Neitzke, G. de Souza, and R. Torres, Microstructural and tribological characterization of dlc coatings deposited by plasma enhanced techniques on steel substrates, Surface and Coatings Technology 389, 125615 (2020).
- Weiser et al. [1995] P. S. Weiser, S. Prawer, R. R. Manory, A. Hoffman, P. J. Evans, and P. J. Paterson, Chemical vapour deposition of diamond onto steel: the effect of a ti implant layer, Surface and Coatings Technology 71, 167 (1995).
- Chang and Wang [2001] C.-L. Chang and D.-Y. Wang, Microstructure and adhesion characteristics of diamond-like carbon films deposited on steel substrates, Diamond and Related Materials 10, 1528 (2001).
- Guo et al. [2010] H. Guo, Y. Qi, and X. Li, Adhesion at diamond/metal interfaces: A density functional theory study, Journal of Applied Physics 107, 033722 (2010).
- Wang et al. [2021] K. Wang, H. Zhou, K. Zhang, X. Liu, X. Feng, Y. Zhang, G. Chen, and Y. Zheng, Effects of ti interlayer on adhesion property of dlc films: A first principle study, Diamond and Related Materials 111, 108188 (2021).
- Hagarová et al. [2024] M. Hagarová, G. Baranová, M. Heželová, M. Truchlý, M. Vojtko, O. Petruš, and D. Csík, High-temperature mechanical and tribological performance of w-dlc coating with cr interlayer on x40crmov5-1 hot work tool steel, Coatings 14, 10.3390/coatings14080971 (2024).
- Ilic et al. [2019] E. Ilic, A. Pardo, T. Suter, S. Mischler, P. Schmutz, and R. Hauert, A methodology for characterizing the electrochemical stability of dlc coated interlayers and interfaces, Surface and Coatings Technology 375, 402 (2019).
- Bentzon et al. [1994] M. Bentzon, K. Mogensen, J. B. Hansen, C. Barholm-Hansen, C. Triholt, P. Holiday, and S. Eskildsen, Metallic interlayers between steel and diamond-like carbon, Surface and Coatings Technology 68-69, 651 (1994).
- Vetter [2014] J. Vetter, 60years of dlc coatings: Historical highlights and technical review of cathodic arc processes to synthesize various dlc types, and their evolution for industrial applications, Surface and Coatings Technology 257, 213 (2014).
- Grill [1999] A. Grill, Diamond-like carbon: state of the art, Diamond and Related Materials 8, 428 (1999).
- Gershman et al. [2022] I. Gershman, A. Mironov, A. Mezrin, E. Torskaya, T. Kuznetsova, V. Lapitskaya, and A. Rogachev, Effect of sp3–sp2 transformation on the wear rate of the dlc coating, Lubricants 10, 10.3390/lubricants10050085 (2022).
- Yamamoto et al. [2012] S. Yamamoto, A. Kawana, H. Ichimura, and C. Masuda, Relationship between tribological properties and sp3/sp2 structure of nitrogenated diamond-like carbon deposited by plasma cvd, Surface and Coatings Technology 210, 1 (2012).
- Li et al. [2020] C. Li, L. Huang, and J. Yuan, Effect of sp3 content on adhesion and tribological properties of non-hydrogenated dlc films, Materials 13, 10.3390/ma13081911 (2020).
- Cheng et al. [2024] C. Cheng, X. Sun, X. He, E. Xu, Q. Wang, R. Cao, W. Shen, F. Dong, Y. Guo, Z. Zhang, and S. Liu, Atomic and electronic properties of the metal/diamond (100) interfaces by first-principles calculations, Surfaces and Interfaces 52, 104916 (2024).
- Ichibha et al. [2018] T. Ichibha, K. Hongo, I. Motochi, N. Makau, G. Amolo, and R. Maezono, Adhesion of electrodes on diamond (111) surface: A dft study, Diamond and Related Materials 81, 168 (2018).
- Wang and Smith [2001] X.-G. Wang and J. R. Smith, Copper/diamond adhesion and hydrogen termination, Phys. Rev. Lett. 87, 186103 (2001).
- Zhang et al. [2022] S. Zhang, L. Rao, W. Shao, Q. He, X. Xing, Y. Zhou, and Q. Yang, Electronic, tensile, and sliding characteristics of the c/ti interface: A first-principles study, Langmuir 38, 15113 (2022).
- Zur and McGill [1984] A. Zur and T. C. McGill, Lattice match: An application to heteroepitaxy, Journal of Applied Physics 55, 378 (1984).
- Pierre et al. [2014] M. D. L. Pierre, M. Bruno, C. Manfredotti, F. Nestola, M. Prencipe, and C. Manfredotti, The (100), (111) and (110) surfaces of diamond: an ab initio b3lyp study, Molecular Physics 112, 1030 (2014).
- Stekolnikov et al. [2002] A. A. Stekolnikov, J. Furthmüller, and F. Bechstedt, Absolute surface energies of group-iv semiconductors: Dependence on orientation and reconstruction, Phys. Rev. B 65, 115318 (2002).
@@
- Bechstedt [2003] F. Bechstedt, Principles of Surface Physics, 1st ed., Advanced Texts in Physics (Springer, Berlin, Heidelberg, 2003).
- Losi et al. [2023] G. Losi, O. Chehaimi, and M. C. Righi, Tribchem: A software for the first-principles, high-throughput study of solid interfaces and their tribological properties, Journal of Chemical Theory and Computation 19, 5231 (2023).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Ab initio molecular dynamics for liquid metals, Phys. Rev. B 47, 558 (1993).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Jain et al. [2015] A. Jain, S. P. Ong, W. Chen, B. Medasani, X. Qu, M. Kocher, M. Brafman, G. Petretto, G.-M. Rignanese, G. Hautier, D. Gunter, and K. A. Persson, Fireworks: a dynamic workflow system designed for high-throughput applications, Concurrency and Computation: Practice and Experience 27, 5037 (2015).
- Mathew et al. [2017] K. Mathew, J. H. Montoya, A. Faghaninia, S. Dwarakanath, M. Aykol, H. Tang, I. heng Chu, T. Smidt, B. Bocklund, M. Horton, J. Dagdelen, B. Wood, Z.-K. Liu, J. Neaton, S. P. Ong, K. Persson, and A. Jain, Atomate: A high-level interface to generate, execute, and analyze computational materials science workflows, Computational Materials Science 139, 140 (2017).
- [42] Custodian, https://github.com/materialsproject/custodian.
- Ong et al. [2013] S. P. Ong, W. D. Richards, A. Jain, G. Hautier, M. Kocher, S. Cholia, D. Gunter, V. L. Chevrier, K. A. Persson, and G. Ceder, Python materials genomics (pymatgen): A robust, open-source python library for materials analysis, Computational Materials Science 68, 314 (2013).
- Mathew et al. [2016] K. Mathew, A. K. Singh, J. J. Gabriel, K. Choudhary, S. B. Sinnott, A. V. Davydov, F. Tavazza, and R. G. Hennig, Mpinterfaces: A materials project based python tool for high-throughput computational screening of interfacial systems, Computational Materials Science 122, 183 (2016).
- Wolloch et al. [2019] M. Wolloch, G. Losi, M. Ferrario, and M. C. Righi, High-throughput screening of the static friction and ideal cleavage strength of solid interfaces, Scientific Reports 9, https://doi.org/10.1038/s41598-019-49907-2 (2019).
- Restuccia et al. [2023] P. Restuccia, G. Losi, O. Chehaimi, M. Marsili, and M. C. Righi, High-throughput first-principles prediction of interfacial adhesion energies in metal-on-metal contacts, ACS Applied Materials & Interfaces 15, 19624 (2023).
- Perdew et al. [1996] J. P. Perdew et al., Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The materials project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Birch [1947] F. Birch, Finite elastic strain of cubic crystals, Phys. Rev. 71, 809 (1947).
- Damiani et al. [2024] E. Damiani, M. Marsili, and M. C. Righi, Tuning the adhesion of diamond/copper interfaces through surface chemical modifications and reconstruction, Carbon 230, 119555 (2024).
- Israelachvili [2011] J. N. Israelachvili, in Intermolecular and Surface Forces, edited by J. N. Israelachvili (Academic Press, San Diego, 2011) 3rd ed., pp. 191–204.
- Young and Goodwin [2011] C. A. Young and A. L. Goodwin, Applications of pair distribution function methods to contemporary problems in materials chemistry, J. Mater. Chem. 21, 6464 (2011).
@@
- Billinge [2008] S. J. Billinge, Nanoscale structural order from the atomic pair distribution function (pdf): There’s plenty of room in the middle, Journal of Solid State Chemistry 181, 1695 (2008).
@@
- Fernandez et al. [2013] M. Fernandez, N. R. Trefiak, and T. K. Woo, Atomic property weighted radial distribution functions descriptors of metal–organic frameworks for the prediction of gas uptake capacity, The Journal of Physical Chemistry C 117, 14095 (2013).
- Kajita and Righi [2016] S. Kajita and M. Righi, A fundamental mechanism for carbon-film lubricity identified by means of ab initio molecular dynamics, Carbon 103, 193 (2016).
- Pacini et al. [2024] A. Pacini, M. Ferrario, S. Loehle, and M. Righi, dvancing tribological simulations of carbon-based lubricants with active learning and machine learning molecular dynamics, Eur. Phys. J. Plus 139, https://doi.org/10.1140/epjp/s13360-024-05348-z (2024).
- Marsili et al. [2005] M. Marsili, O. Pulci, F. Bechstedt, and R. Del Sole, Electronic structure of the surface: Solution by self-consistent many-body calculations, Phys. Rev. B 72, 115415 (2005).

### v2

@@ 3.5 Charge density redistribution at the interface
The structural modifications of the diamond reconstructions directly affect the electronic charge redistribution at the interface. Previous studies have shown that the redistribution of electronic charge upon contact, quantified by the descriptor (Eq. 2), correlates well with adhesion across a wide variety of homogeneous interfaces, from weakly interacting van der Waals systems (e.g., graphene) to covalently bonded materials such as diamond [56], as well as heterogeneous metallic junctions [46].
In the present case, the semiconductor-metal interaction introduces additional complexity, as surface carbon chains can undergo rehybridization and other structural rearrangements, modifying the local electronic environment. These factors, together with the broader variability of bond character, naturally increase the scatter. Nevertheless, a good correlation between adhesion energy and is still present, with (Fig. 9), supporting as a reliable predictor also for this class of interfaces. Notably, even for homogeneous interfaces the correlation is not perfect: in the reference study the metallic class (the most populated subset) corresponds to within a linear model, reflecting the sensitivity of to the detailed bonding chemistry at the interface [56].
To gain insight into the origin of the residual scatter, here we analyze two opposite outliers: C(111)-rec/Al exhibits an anomalously large relative to its modest adhesion, whereas C(111)/Ti shows comparatively low despite high adhesion. To rationalize these systems, we inspected the planar averaged charge displacement , which reveals that the spatial extent of the electronic response differs markedly between the two interfaces. In C(111)-rec/Al (Fig.10(a)), is not confined to the immediate contact region but extends deep into the Al slab. This is consistent with metallic screening: delocalized electrons redistribute over several atomic layers, forming an interface dipole and yielding an oscillatory on the Al side. This screening can lead to a large (defined as the integral of within the interfacial window) without translating into a proportional adhesion gain, because it is less directly tied to the formation of short, directional interfacial bonds. Consistently, the average interfacial separation is relatively large (C–Al = 2.14 Å), and the Al slab shows negligible first interlayer relaxation (: from 2.334 to 2.336 Å), both indicative of limited chemisorption-driven accommodation.
In contrast, for C(111)/Ti (Fig.10(b)) the perturbation extends into the diamond slab, indicating that the interaction modifies the carbon electronic structure beyond the topmost layer. This behavior is compatible with strong chemisorption (consistent with C 2–Ti 3 hybridization) that enhances adhesion, and it is supported by the shorter interfacial distance (C–Ti = 1.95 Å) and by a clear substrate relaxation on the Ti side (: from 2.174 to 2.262 Å, expansion). In this regime, adhesion can be large even if computed within a fixed interfacial region is comparatively low, because part of the charge rearrangement is spatially distributed into the carbon slab rather than being concentrated only at the immediate contact. Overall, these two outliers emphasize that quantifies the magnitude of charge rearrangement within a chosen interfacial region, but the adhesion also depends on whether that rearrangement reflects long-range metallic screening (as in Al) or efficient bond-forming chemisorption with structural accommodation (as in Ti).
4 CONCLUSIONS
@@
In this work, a high-throughput ab initio investigation of diamond/metal interfaces has been carried out, systematically analyzing the adhesion of 60 combinations involving a broad range of metallic substrates and four low-index diamond terminations (C(111), C(111)-rec, C(110), C(100)), representative of the different types of bonds in DLC coatings. The study employs an automated workflow implemented within the TribChem software, which ensures a consistent and reproducible protocol for the construction and characterization of solid interfaces. In particular, the interface supercells were built using the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while minimizing both lattice mismatch and angular distortion. All interfaces were constructed with a metallic strain below 3, ensuring that the calculated adhesion energies are not affected by unphysical distortions.
@@
The results show that adhesion is primarily governed by the intrinsic reactivity of the metallic component and by the type of carbon hybridization present at the surface of the diamond. Carbide-forming metals such as Cr, Ti, W and V exhibit the strongest adhesion, in agreement with their experimentally recognized role as interlayer materials that enhance DLC adhesion to metallic substrates. In contrast, noble and weakly reactive metals (Ag, Au, Cu, Zn, Mg) display poor adhesion due to their filled valence shells and low surface energies. Among the diamond terminations, the C(111) surface with sp3 bonding shows the highest adhesion, whereas surface graphitization, as observed in C(111)-rec, considerably weakens interfacial bonding.
@@
By comparing the work of separation () of the diamond/metal interface with the cohesive energy of the corresponding bulk metal, it is possible to predict the likely fracture location under tensile load. Interfaces involving C(111)-rec consistently exhibit fracture at the interface, reflecting the weak bonding due to sp2 terminations. For metals with low cohesive energies, such as Mg, Zn and Al, fracture can occur within the metal bulk when in contact with highly reactive diamond surfaces, indicating that the interface is stronger than the metallic cohesion. Conversely, transition metals with partially filled d-shells (e.g., W, Cr, Mo, Fe) exhibit high bulk cohesion, preventing failure from taking place within the metal phase.
Analysis of the radial distribution function provides insights on the structural response of diamond surface reconstructions upon contact with metals. In C(111)-rec, the peak at approximately 1.44 Å, corresponding to the zig-zag surface chains, is largely preserved in the presence of weakly interacting metals such as Au and Ag, indicating minimal perturbation of the sp2 network. In contrast, Cr, Al, Cu and Pt induce significant dimerization and partial rehybridization of the surface sp2 bonds. In general, reactive metals like W, Cr, Mo, Fe, V and Ti are able to reduce the number of surface sp2 bonds by up to 50. Hence, it can be concluded that strong perturbation and reduction in the number of surface sp2 bonds in diamond coatings directly enhance adhesion, explaining why these metals are commonly used as interlayers to improve bonding between DLC films and metallic substrates.
The structural modifications are mirrored in an electronic rearrangement at the interface. Volumetric charge density differences show that metals inducing stronger structural perturbations promote more pronounced charge rearrangement, while weaker interacting metals cause only minor redistribution. Quantitatively, the charge redistribution descriptor, , remains strongly correlated with adhesion, confirming that electronic charge redistribution upon interface formation is a good predictor of adhesion strength, even in the presence of semiconductor-metal bonding.
In summary, by focusing on ideal crystalline interfaces, this work aimed at establishing physically grounded and broadly applicable correlations linking adhesion to intrinsic structural and electronic descriptors, without resorting to system specific models and by exploiting the accuracy of first principles calculations. In the future, machine learning approaches could leverage these insights to address more realistic defective or amorphous interfaces and larger length/time scales.
@@ Acknowledgements.
These results are part of the ”Advancing Solid Interface and Lubricants by First Principles Material Design (SLIDE)” project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 865633). The authors acknowledge funding by the European Union - NextGenerationEU (National Sustainable Mobility Center CN00000023, Italian Ministry of University and Research Decree n. 1033 - 17/06/2022, Spoke 11 - Innovative Materials Lightweighting) and by ICSC-Centro Nazionale di Ricerca in High Performance Computing,Big Data and Quantum Computing, funded by EuropeanUnion-NextGenerationEU. The opinions expressed are those of the authors only and should not be considered as representative of the European Union or the European Commission’s official position. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support.References
@@
- [1] (2004) Materials science and fabrication processes for a new MEMS technology based on ultrananocrystalline diamond thin films. Journal of Physics: Condensed Matter 16 (16), pp. R539. External Links: Document Cited by: §1.
- [2] (2009) Tribocorrosion behaviour of DLC-coated 316L stainless steel. Wear 267 (5), pp. 860–866. External Links: Document Cited by: §1.
@@
- [3] (2003) Principles of surface physics. 1 edition, Advanced Texts in Physics, Springer, Berlin, Heidelberg. External Links: ISBN 978-3-540-00635-0, Document, ISSN 1439-2674 Cited by: §1.
- [4] (1994) Metallic interlayers between steel and diamond-like carbon. Surface and Coatings Technology 68-69, pp. 651–655. External Links: Document Cited by: §1.
@@
- [5] (2008) Nanoscale structural order from the atomic pair distribution function (PDF): there’s plenty of room in the middle. Journal of Solid State Chemistry 181 (7), pp. 1695–1700. External Links: Document Cited by: §3.4.
- [6] (1947) Finite elastic strain of cubic crystals. Phys. Rev. 71, pp. 809–824. External Links: Document Cited by: §2.
- [7] (2001) Microstructure and adhesion characteristics of diamond-like carbon films deposited on steel substrates. Diamond and Related Materials 10 (8), pp. 1528–1534. External Links: Document Cited by: §1.
- [8] (2024) Atomic and electronic properties of the metal/diamond (100) interfaces by first-principles calculations. Surfaces and Interfaces 52, pp. 104916. External Links: Document Cited by: §1.
- [9] Custodian. Note: https://github.com/materialsproject/custodian Cited by: §2.
- [10] (2024) Tuning the adhesion of diamond/copper interfaces through surface chemical modifications and reconstruction. Carbon 230, pp. 119555. External Links: Document Cited by: §1, §3.1.
- [11] (2014) The (100), (111) and (110) surfaces of diamond: an ab initio B3LYP study. Molecular Physics 112 (7), pp. 1030–1039. External Links: Document Cited by: §1.
- [12] (2005) Biomedical applications of diamond-like carbon (DLC) coatings: a review. Surface and Coatings Technology 200 (7), pp. 2518–2524. External Links: Document Cited by: §1.
@@
- [13] (2013) Atomic property weighted radial distribution functions descriptors of metal–organic frameworks for the prediction of gas uptake capacity. The Journal of Physical Chemistry C 117 (27), pp. 14095–14105. External Links: Document Cited by: §3.4.
- [14] (2022) Effect of sp3–sp2 transformation on the wear rate of the DLC coating. Lubricants 10 (5), pp. 85. External Links: Document Cited by: §1.
- [15] (1999) Diamond-like carbon: state of the art. Diamond and Related Materials 8 (2), pp. 428–434. External Links: Document Cited by: §1.
- [16] (2010) Adhesion at diamond/metal interfaces: a density functional theory study. Journal of Applied Physics 107 (3), pp. 033722. External Links: Document Cited by: §1, §1, §3.1.
- [17] (2024) High-temperature mechanical and tribological performance of W-DLC coating with Cr interlayer on X40CrMoV5-1 hot work tool steel. Coatings 14 (8), pp. 971. External Links: Document Cited by: §1.
- [18] (2003) A review of modified DLC coatings for biological applications. Diamond and Related Materials 12 (3), pp. 583–589. External Links: Document Cited by: §1.
- [19] (2018) Adhesion of electrodes on diamond (111) surface: a dft study. Diamond and Related Materials 81, pp. 168–175. External Links: Document Cited by: §1.
- [20] (2019) A methodology for characterizing the electrochemical stability of DLC coated interlayers and interfaces. Surface and Coatings Technology 375, pp. 402–413. External Links: Document Cited by: §1.
- [21] (2011) In Intermolecular and Surface Forces, J. N. Israelachvili (Ed.), pp. 191–204. Cited by: §3.2.
- [22] (2015) FireWorks: a dynamic workflow system designed for high-throughput applications. Concurrency and Computation: Practice and Experience 27 (17), pp. 5037–5059. External Links: Document Cited by: §2.
- [23] (2013) Commentary: the materials project: a materials genome approach to accelerating materials innovation. APL Materials 1 (1), pp. 011002. External Links: Document Cited by: §2, §3.1.